OpenStudy (anonymous):

An arch has the shape of the parabola y = 36 − x2. A rectangle is fit under the arch by selecting a point (x, y) on the parabola. (a) Express the area A of the rectangle in terms of x. (b) If x = 1, the rectangle has base 2 and height 35. Find the base of a second rectangle that has the same area.

OpenStudy (anonymous):

thats fairly easy too

OpenStudy (anonymous):

not the best picture , but good enough

OpenStudy (anonymous):

assuming its symmetric about y axis , which it most likely is the area is (2x) y , ie base times height

OpenStudy (anonymous):

A=2xy but y= 36-x^2 so A=2x ( 36-x^2)

OpenStudy (anonymous):

so base is 2x and all theyare asking for second part is to solve A= 70

OpenStudy (anonymous):

72x -2x^3 =70

OpenStudy (anonymous):

this is a cubic, requires polynomial division etc like all the others you have been doing

OpenStudy (anonymous):

Okay got it! You're amazing!